OpenStudy (anonymous):

if f(x)=sqroot (4-(x^2)) then the domain is (-infinity,4] right? how can I get the range without sketching?

OpenStudy (anonymous):

First the domain is not (- infinity 4). It is [-2,2].

OpenStudy (anonymous):

The range is [0,2]. You can say that since x lies in [-2,2] and the function returns same value for x and -x. Minimum is at x=2. and max at x=0

OpenStudy (anonymous):

you got the domain wrong. the domain of sqroot is [0,+inf) thus 4-(x^2)>=0 so x^2 <= 4 so \[x \le \pm 2\]

OpenStudy (anonymous):

sqroot is continuous function so if you find min and max of 4-x^2 while considering the domain and sqroot it - you will find your range

OpenStudy (anonymous):

oh yes you are right i got the domain from x^2<= 4 thanks a lot not x

OpenStudy (anonymous):

the min/max of the function in sqroot are 0 and 4 respectively thus your range is [0,2]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

something to note about this type of questions is that sketching helps. if you draw a quick sketch of the y=-x^2+4 parabula ("sad" parabula because of the negative sign in front of x^2 with its peak at (0,4) )

OpenStudy (anonymous):

you are 100% right visual calculus is very helpful , thanks for the tip really appreciate it